A researcher examined a random sample of 300 homes in a small city and found that 52 had solar panels installed on their roofs. Use the sample to construct a 99% confidence interval for the proportion of all homes in the city that have solar panels installed on their roofs. What is the margin of error for the 99% confidence interval? Round to three decimal places.
Answer #1

Solution :

Given that,

n = 300

x = 52

Point estimate = sample proportion = \hat p = x / n = 52/300 = 0.173

1 - \hat p = 1-0.173 = 0.827

At 99% confidence level

\alpha = 1-0.99% =1-0.99 =0.01

\alpha/2 =0.01/ 2= 0.005

Z\alpha/2 = Z0.005 = 2.576

Z\alpha/2 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt(0.173*(0.827) /300 )

= 0.056

A 99% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.173-0.056 < p < 0.173+0.056

0.117 < p < 0.229

(0.117 ,0229 )